OpenStudy (hyuna301):

Solve : arctan {1/(cos2@)^1/2} - arctan {(cos2@)^1/2}

OpenStudy (hyuna301):

@myininaya

myininaya (myininaya):

I'm going to use x instead of @ does solve mean simplify because I see no equation to actually solve \[\arctan(\frac{1}{(\cos(2x))^\frac{1}{2}})- \arctan((\cos(2x))^\frac{1}{2})\]

OpenStudy (hyuna301):

Actually, u = given expression and I need to find 'sin u'

myininaya (myininaya):

oh so we want to use difference identity for sine

myininaya (myininaya):

\[\sin(g-h)=\sin(g)\cos(h)-\sin(h) \cos(g) \\ \text{ where } g=\arctan(\frac{1}{(\cos(2x))^\frac{1}{2}}) \\ \text{ and } h=\arctan((\cos(2x))^\frac{1}{2})\]

myininaya (myininaya):

we can write sin(g) , cos(h), sin(h), and cos(g) as algebraic expressions

myininaya (myininaya):

let me know if you have questions so far I will give you an example of the algebraic expression if you need

OpenStudy (hyuna301):

I was unable to find the hypotaneous

myininaya (myininaya):

for the algebraic expression part?

myininaya (myininaya):

let's look at sin(g)....

myininaya (myininaya):

when I say sin(g) this is short for \[\sin(\arctan(\frac{1}{(\cos(2x))^\frac{1}{2}}))\]

OpenStudy (hyuna301):

Yeah .. I simplified the given expression and got tan u but was unable to get sin u

myininaya (myininaya):

\[g=\arctan(\frac{1}{(\cos(2x))^\frac{1}{2}}) \\ \tan(g)=\frac{1}{(\cos(2x))^\frac{1}{2}} =\frac{opp}{adj}\]

myininaya (myininaya):